a) Find the volume of CO2 formed at 25 °C and 103 kPa when 0.316 g sodium carbonate is added to 12.0 mL of 0.110 M phosphoric acid.

3 Na2CO3 (s) + 2 H3PO4 (aq) −> 2 Na3PO4 (aq) + 3 H2O (l) + 3 CO2 (g)


b) If the leftover amount of excess reactant is thermally decomposed, what volume of gas will be formed at the same conditions as part a)? Show the chemical equation for the decomposition reaction in your answer.


1
Expert's answer
2022-01-20T11:00:02-0500

Molar mass of NaCO3 = 105.9

0.316/105.9=0.003

PV=nRT

103×V= 0.113×4.185×298

103V= 3.742

= 0.036ml